a teacher is four times as old as a student. in 30 years, the students age will be half of the teachers age. How old are they now?
Answer by lwsshak3(11628)

a teacher is four times as old as a student. in 30 years, the students age will be half of the teachers age. How old are they now
..
let x=student's current age
4x=teacher's current age
30 years hence
student's age=x+30
teacher's age=4x+30
(x+30)/(4x+30)=1/2
2x+60=4x+30
2x=30
x=15
ans:
student's current age=15
teacher's current age=60